OpenStudy (megannicole51):

Does the series converge or diverge? N=1 and b=infinity of (1/2)^n+(2/3)^n

OpenStudy (psymon):

Well, you have the addition of two geo series really. \[\sum_{n=1}^{\infty}(\frac{ 1 }{ 2 })^{n} + \sum_{n=1}^{\infty}(\frac{ 2 }{ 3 })^{n}\]

OpenStudy (megannicole51):

Yup:)

OpenStudy (psymon):

For a geo-series in the form of ar^n, if the absolute value of r, (1/2) and (2/3) would be your r's in this case, is between 0 and 1, the sequence diverges. What it diverges to is \[\frac{ a }{ 1-r } \]But then again, if you dont need the actual sum I suppose you could stop xD

OpenStudy (anonymous):

\[\{\left(\frac{1}{2}\right)^n+\left(\frac{2}{3}\right)^n\}_{n=1}^{\infty}\] a sequence, or a series ?

OpenStudy (anonymous):

oh nvm it says series what @Psymon wrote

OpenStudy (anonymous):

except where @Psymon wrote "diverge" it should be "converge"

OpenStudy (psymon):

Yeah, accident xD

OpenStudy (anonymous):

no doubt a typo "For a geo-series in the form of ar^n, if the absolute value of r, (1/2) and (2/3) would be your r's in this case, is between 0 and 1, the sequence CONverges. What it CONerges to is "

OpenStudy (psymon):

guess I was being lazy about the specifics of it x_X

OpenStudy (megannicole51):

Oh thank you!!

OpenStudy (megannicole51):

Brb!

OpenStudy (psymon):

Lol, okay O.o

OpenStudy (megannicole51):

back:D sorry i had to get gas for my car!

OpenStudy (anonymous):

was it expensive?

OpenStudy (megannicole51):

yeah as always:(

OpenStudy (anonymous):

lol did you add this stuff up yet, or are you done?

OpenStudy (megannicole51):

its actually pretty cheap though but it still sucks ya know lol

OpenStudy (megannicole51):

no i havent added anything yet

OpenStudy (megannicole51):

so do i just put 1 in for n?

OpenStudy (psymon):

Least ya got a car, lol.

OpenStudy (anonymous):

no you do as before \[\sum_{n=1}^{\infty}\frac{1}{2^n}=\frac{\frac{1}{2}}{1-\frac{1}{2}}=1\]

OpenStudy (megannicole51):

true!

OpenStudy (megannicole51):

im very thankful

OpenStudy (anonymous):

mine likes lots of gas

OpenStudy (psymon):

Doesnt matter for me, I couldnt drive even if I had a car xD

OpenStudy (megannicole51):

where did u get (1/2^n)?

OpenStudy (anonymous):

\(\left(\frac{1}{2}\right)^n=\frac{1}{2^n}\) i just wrote it differently

OpenStudy (anonymous):

i forget whether you had to add this up, or just say if it converged or not

OpenStudy (megannicole51):

do i have to do (2/3)^n as well?

OpenStudy (megannicole51):

and idk what u mean? like add up both sides?

OpenStudy (anonymous):

yes the same way \[\sum_{n=1}^{\infty}\left(\frac{2}{3}\right)^n=\frac{\frac{2}{3}}{1-\frac{2}{3}}\]

OpenStudy (anonymous):

because \[\sum_{n=1}^{\infty}(\frac{1}{2})^n+(\frac{2}{3})^n=\sum_{n=1}^{\infty}(\frac{1}{2})^n+\sum_{n=1}^{\infty}(\frac{2}{3})^n\]so you can add them up separately

OpenStudy (watchmath):

Just to add a small think. If the objective is just to find out if this series is convergent, we don't really need to find the sum. After we split the series as psymon did, we say that the series is the sum of two geometric series each with ratio less than 1. So each series is convergent and so does the sum.

OpenStudy (megannicole51):

@satellite73 so if u add them up its 3? @watchmath thank you...so its converges as long as the limits aren't zero right?

OpenStudy (watchmath):

no, we don't look at any limit here. We use the fact that a geometric series is convergent if only if the ratio is less than 1.

OpenStudy (anonymous):

no if the limit of the sum was somehow zero, then it would converge to zero and yes, the sum is 3

OpenStudy (megannicole51):

so the fact that the sum is 3 tells me what exactly? im sorry im just getting a little mixed up...theres so much information that goes into these problems!

OpenStudy (anonymous):

the fact that you can actually add these up means the sum converges, it converges to that number

OpenStudy (megannicole51):

awesome:) makes sense to me!

OpenStudy (anonymous):

but you might have a sum that you know converges but you could not be able to find that number you would just know it converged by some test or other, but you would not know how to actually compute the sum

OpenStudy (anonymous):

in this case you not only know that it converges, but you know what it converges to, namely 3

OpenStudy (megannicole51):

perfect that makes sense!

OpenStudy (megannicole51):

thank you!

OpenStudy (anonymous):

yw

OpenStudy (megannicole51):

do u mind helping me with a couple more? ill repost them

OpenStudy (anonymous):

if i can

OpenStudy (anonymous):

post in a new thread, you will get more responses some may actually be right